In 1993, Crane Company completed the construction of a building at a cost of $2,020,000 and first occupied it in January 1994. It was estimated that the building will have a useful life of 40 years and a salvage value of $60,000 at the end of that time.

Early in 2004, an addition to the building was constructed at a cost of $505,000. At that time, it was estimated that the remaining life of the building would be, as originally estimated, an additional 30 years, and that the addition would have a life of 30 years and a salvage value of $20,200.

In 2022, it is determined that the probable life of the building and addition will extend to the end of 2053, or 20 years beyond the original estimate.

Compute the annual depreciation to be charged, beginning with 2022. (Round answer to 0 decimal places, e.g. 45,892.)
Answer #1
Annual depreciation to be charged, beginning with 2022 24435
Workings:
Cost 2020000
Less: Salvage value 60000
Depreciable cost 1960000
Useful life 40
Annual depreciation from 1994 through 2003 49000
Cost 2020000
Add: Additions 505000
Total cost 2525000
Less: Accumulated depreciation from 1994 through 2003 490000 =49000*10
Book value, Jan 2004 2035000
Less: Salvage value 80200 =60000+20200
Depreciable cost 1954800
Useful life 30
Annual depreciation from 2004 through 2021 65160
Book value, Jan 2004 2035000
Less: Accumulated depreciation from 2004 through 2021 1172880 =65160*18
Book value, Jan 2022 862120
Less: Salvage value 80200 =60000+20200
Depreciable cost 781920
Remaining Useful life 32 =12+20
Annual depreciation expense—building 24435